# In-SRAM Radiant Foam Rendering on a Graph Processor

### v1

@@ Abstract
We present a fully in-SRAM, distributed renderer for the Radiant Foam Voronoi-cell volumetric representation on the Graphcore Mk2 IPU, a many-core accelerator with tile-local SRAM and explicit inter-tile communication. Our system shards the scene across tiles and forwards rays between shards through a hierarchical routing overlay, enabling ray marching entirely from on-chip SRAM with predictable communication. On Mip-NeRF 360 scenes, the system attains near-interactive throughput ($\approx$1 fps at $640\times 480$) with image and depth quality close to the original GPU-based Radiant Foam implementation, while keeping all scene data and ray state in on-chip SRAM. Beyond demonstrating feasibility, we analyze routing, memory, and scheduling bottlenecks that inform how future distributed-memory accelerators can better support irregular, data-movement-heavy rendering workloads.
@@ 1 Introduction
We therefore choose a volumetric representation whose ray traversal is local in the representation’s adjacency graph. Specifically, we adopt Radiant Foam [10]: because its Voronoi cells form a connected adjacency structure. Rendering a single ray can be viewed as walking the Voronoi diagram: the ray starts in the cell containing the camera position and then advances cell by cell by checking only the current cell and its immediate neighbors. To execute this ray traversal on an IPU, we partition Voronoi cells into SRAM-sized shards and place each shard on a tile (Fig. 1). Whenever a ray moves into a Voronoi cell whose shard is owned by a different tile, its state must be transferred to that tile; similarly, once a ray completes, its pixel contribution need to be sent to the tile that owns that region of the output image. Rather than allowing arbitrary all-to-all messaging, we structure these transfers through a static hierarchical routing overlay so that each tile communicates through a small, predictable set of links(Fig. 1).
@@ 2.1 Radiance-field methods
Complementary to these algorithmic speedups on GPUs, a growing line of work explores custom hardware accelerators for neural rendering. For NeRF, several FPGA- and ASIC-based designs co-design sampling, encoding, and MLP datapaths to reduce latency and power on a single device [29, 35, 48, 51, 26, 38]. For 3D Gaussian Splatting, recent processors and architectures specialize sorting, rasterisation, and memory hierarchies for Gaussian primitives [27, 25, 50, 14, 49, 40]. These accelerators demonstrate that radiance-field methods benefit from hardware specialization, but they typically entail long design and verification cycles and are tightly coupled to a particular rendering pipeline. Adapting them to new radiance-field architectures or training regimes often requires substantial re-engineering, whereas programmable many-core processors with distributed SRAM, as we explore in this paper with an IPU, can support a wider range of designs via software while still enforcing strict communication locality.
@@ 2.3 Emerging AI accelerators beyond GPUs
Alongside GPUs and application-specific accelerators, several families of AI accelerators now target large-scale deep learning with different architectures. Wafer-scale engines such as the Cerebras WSE-2 integrate hundreds of thousands of small compute cores and tens of gigabytes of on-chip SRAM on a single wafer, connected by a mesh-like on-chip network [30]. Such chips are best viewed as distributed-memory machines, where each core owns a small local store and communication follows explicit routes over the on-wafer fabric [13]. Spatial dataflow accelerators such as Google’s Tensor Processing Units (TPUs) organise computation around systolic matrices attached to high-bandwidth off-chip memory, exposing coarse-grained array-style parallelism rather than the fine-grained thread hierarchy of a GPU [20]. Finally, many-core ML processors such as the Graphcore IPU [4, 11] (see Sec. 3.2) populate the chip with over a thousand tiles, each with its own SRAM and statically scheduled communication routes.
@@ 3.2 Graphcore IPU
Programs are expressed in Graphcore’s Poplar framework as kernels mapped to tiles and operating on explicitly placed tensors [11, 4]. Execution proceeds in bulk-synchronous phases: in each phase, tiles (i) compute on their local data, (ii) reach a global synchronisation point, and then (iii) exchange the next slice of data along pre-defined routes. Communication between tiles is scheduled at compile time, so exchange routes and buffer sizes are fixed ahead of execution, requiring bounded communication degrees and static message patterns. Within a tile, each kernel is executed by a small group of worker threads that share tile-local SRAM. So, concurrent updates must be organized so that each worker writes to disjoint memory regions between synchronization points.
@@ 4.1 System overview
We organize the IPU tiles into three roles (Fig. 1). Ray tracer tiles each store a shard of the Radiant Foam scene and a corresponding slice of the framebuffer, and perform ray marching on rays traversing their local Voronoi cells. Router tiles forward rays to its next destination. A single ray generator tile creates new rays from the current camera parameters. All tiles execute in parallel on the IPU: tracers advance rays through local data, routers forward rays toward their next destination, and the generator injects new rays. Periodically, the host reads back the distributed framebuffer slices and assembles the final image.
@@ 4.4 Ray representation
Table 1: Per-ray payload variants.
@@ 4.6 Ray generator and batched injection
(a) Ground truth
@@
Figure 8: RGB reconstruction on the Mip-NeRF360 Garden scene. Our in-SRAM IPU renderer closely matches the Radiant Foam reference while keeping the entire model and all ray state on chip.
@@ 4.8 Execution on IPU
The IPU follows a bulk-synchronous parallel (BSP) model: computation proceeds in phases separated by global synchronization points where data exchange occurs. We exploit this by grouping all tiles (ray tracers, routers, and generator) into a single conccurrent compute step that execute parallel, then alternating compute and exchange programs. The resulting IPU program consists of the following subprograms or kernels:
@@ 5.1 Reconstruction quality
Fig. 8 compares our IPU implementation to the Radiant Foam ground truth on the Garden scene. Despite running entirely from on-chip SRAM with simplified per-cell color (constant RGB instead of spherical harmonics), the IPU renderer preserves fine detail with only minor view-dependent effects difference.
@@
Fig. 9 visualises the same comparison: full-, mixed-, and half-precision payloads produce outputs that are visually indistinguishable.
@@
Figure 10: color and depth on the Stump scene. Depth is computed at the 50% transmittance quantile and tracks scene geometry smoothly.
@@ 5.2 Scene partitioning and on-chip memory footprint
Figure 11: KD-tree partitioning of Bicycle into 1024 clusters (scene partitions). Each cluster corresponds to one leaf of the k-d tree and becomes a scene partition on a ray-tracer tile. For each cluster, the four subplots report the number of local cells, neighbor cells, adjacency entries, and neighboring clusters, respectively.
@@ 5.3 Throughput and batching
Fusing all subprograms( Sec. 4.8) into a single device-side loop shifts the majority of runtime from host engine launches and global barriers into tile-local compute-set execution (Detailed comparison can be found inAppendix C, Fig. 20,21). We further accelerate execution by tuning the repeat count $RC$, the number of compute–exchange iterations executed on device before reading back the output image and updating the camera.
@@
Table 3: Row scanning at $640{\times}480$ with a 200-iteration drain after injection.
@@ 5.4 Routing path length and ray lifetime
(b) Histogram of router and tracer hops per pixel until termination.
Figure 14: Bicycle scene path statistics. The hop map (a) reveals where rays repeatedly cross partition boundaries, while the histogram (b) summarizes the distribution of router and tracer hops over all pixels.
@@ 5.5 Router capacity and ray payload precision
After program memory allocation, each router tile is left with roughly $\approx 576$ kB of SRAM to ten buffers (five inputs, five outputs), so each buffer holds about 57.6 kB of rays. Router capacity therefore depends directly on the ray payload size from  Sec. 4.4. Tab. 4 reports the implied per-buffer limits for several layouts.
@@ 5.6 Limitations and artifacts
Figure 16: Ray-origin mismatch artifact in the Playroom scene when the camera pose changes mid-batch. Rays launched under the old pose later reconstruct their origins using the updated pose, causing misaligned directions and results in black or white speckle.
@@ Appendix A Radiant Foam scene representation
Intersecting these half-spaces over all neighbors reconstructs the full Voronoi cell of $p_{c}$ (Figure 17(c)).
Appendix B Scene partitioning: KD-tree vs Octree
@@ B.1 Octree with per-partition point cap $\leq$ 5000 points
Limiting points per partition greatly increases the number of partitions (>>3500), exceeding the available tiles and making a one-tile-per-partition mapping infeasible.
@@
B.2 Octree with fixed partition count (1024 partitions).
Forcing 1024 partitions yields highly unbalanced partitions: some partitions exceed $2\times 10^{4}$ local cells and $3\times 10^{5}$ adjacency entries, surpassing 624 kB SRAM per tile even before queues and workspaces.
@@ D.1 Bicycle scene
(a) Rendered RGB view.
@@ Appendix F Additional qualitative results
Figure 25 and 26 present paired color and depth outputs for three scenes each. Depths are computed from the 50% transmittance quantile, yielding smooth, geometry-aligned maps that corroborate the radiance-field traversal and accumulation behaviour observed in the main results.
(a) Bicycle RGB
(b) Bicycle depth
(c) Garden RGB
(d) Garden depth
(e) Stump RGB
(f) Stump depth
Figure 25: color and depth for Bicycle, Garden, and Stump scenes
(a) Bonsai RGB
(b) Bonsai depth
(c) Kitchen RGB
(d) Kitchen depth
(e) Counter RGB
(f) Counter depth
Figure 26: color and depth for Bonsai, Kitchen, and Counter scenes

### v2

# Radiant Foam Rendering on a Graph Processor

@@ Abstract
We present a fully in-SRAM, distributed renderer for the Radiant Foam Voronoi-cell volumetric representation on the Graphcore Mk2 IPU (Intelligent Processing Unit), a many-core accelerator with tile-local SRAM and explicit inter-tile communication. Our system shards the scene across tiles and forwards rays between shards through a hierarchical routing overlay, enabling ray marching entirely from on-chip SRAM with predictable communication. On Mip-NeRF 360 scenes, the system attains near-interactive throughput ($\approx$1 fps at $640\times 480$) with image and depth quality close to the original GPU-based Radiant Foam implementation, while keeping all scene data and ray state in on-chip SRAM. Beyond demonstrating feasibility, we analyze routing, memory, and scheduling bottlenecks that inform how future distributed-memory accelerators can better support irregular, data-movement-heavy rendering workloads.
@@ 1 Introduction
We therefore choose a volumetric representation whose ray traversal is local in the representation’s adjacency graph. Specifically, we adopt Radiant Foam [10]: because its Voronoi cells form a connected adjacency structure. Rendering a single ray can be viewed as walking the Voronoi diagram: the ray starts in the cell containing the camera position and then advances cell by cell by checking only the current cell and its immediate neighbors. To execute this ray traversal on an IPU, we partition Voronoi cells into SRAM-sized shards and place each shard on a tile (Fig. 1). Whenever a ray moves into a Voronoi cell whose shard is owned by a different tile, its state must be transferred to that tile; similarly, once a ray completes, its pixel contribution needs to be sent to the tile that owns that region of the output image. Rather than allowing arbitrary all-to-all messaging, we structure these transfers through a static hierarchical routing overlay so that each tile communicates through a small, predictable set of links(Fig. 1).
@@ 2.1 Radiance-field methods
Complementary to these algorithmic speedups on GPUs, a growing line of work explores custom hardware accelerators for neural rendering. For NeRF, several FPGA- and ASIC-based designs co-design sampling, encoding, and MLP datapaths to reduce latency and power on a single device [29, 35, 48, 51, 26, 38]. For 3D Gaussian Splatting, recent processors and architectures specialize sorting, rasterization, and memory hierarchies for Gaussian primitives [27, 25, 50, 14, 49, 40]. These accelerators demonstrate that radiance-field methods benefit from hardware specialization, but they typically entail long design and verification cycles and are tightly coupled to a particular rendering pipeline. Adapting them to new radiance-field architectures or training regimes often requires substantial re-engineering, whereas programmable many-core processors with distributed SRAM, as we explore in this paper with an IPU, can support a wider range of designs via software while still enforcing strict communication locality.
@@ 2.3 Emerging AI accelerators beyond GPUs
Alongside GPUs and application-specific accelerators, several families of AI accelerators now target large-scale deep learning with different architectures. Wafer-scale engines such as the Cerebras WSE-2 integrate hundreds of thousands of small compute cores and tens of gigabytes of on-chip SRAM on a single wafer, connected by a mesh-like on-chip network [30]. Such chips are best viewed as distributed-memory machines, where each core owns a small local store and communication follows explicit routes over the on-wafer fabric [13]. Spatial dataflow accelerators such as Google’s Tensor Processing Units (TPUs) organize computation around systolic matrices attached to high-bandwidth off-chip memory, exposing coarse-grained array-style parallelism rather than the fine-grained thread hierarchy of a GPU [20]. Finally, many-core ML processors such as the Graphcore IPU [4, 11] (see Sec. 3.2) populate the chip with over a thousand tiles, each with its own SRAM and statically scheduled communication routes.
@@ 3.2 Graphcore IPU
Programs are expressed in Graphcore’s Poplar framework as kernels mapped to tiles and operating on explicitly placed tensors [11, 4]. Execution proceeds in bulk-synchronous phases: in each phase, tiles (i) compute on their local data, (ii) reach a global synchronization point, and then (iii) exchange the next slice of data along pre-defined routes. Communication between tiles is scheduled at compile time, so exchange routes and buffer sizes are fixed ahead of execution, requiring bounded communication degrees and static message patterns. Within a tile, each kernel is executed by a small group of worker threads that share tile-local SRAM. So, concurrent updates must be organized so that each worker writes to disjoint memory regions between synchronization points.
@@ 4.1 System Overview
We organize the IPU tiles into three roles (Fig. 1). Ray tracer tiles each store a shard of the Radiant Foam scene and a corresponding slice of the framebuffer, and perform ray marching on rays traversing their local Voronoi cells. Router tiles forward rays to their next destination. A single ray generator tile creates new rays from the current camera parameters. All tiles execute in parallel on the IPU: tracers advance rays through local data, routers forward rays toward their next destination, and the generator injects new rays. Periodically, the host reads back the distributed framebuffer slices and assembles the final image.
@@ 4.4 Ray representation
Table 1: Per-ray payload variants used in our renderer.
@@ 4.6 Ray generator and batched injection
(a) Original Radiant Foam rendering results on GPU)
@@
Figure 8: RGB reconstruction on the Mip-NeRF360 Garden scene. Our in-SRAM IPU renderer closely matches the original Radiant Foam implementation while keeping the entire model and all ray state on chip.
@@ 4.8 Execution on IPU
The IPU follows a bulk-synchronous parallel (BSP) model: computation proceeds in phases separated by global synchronization points where data exchange occurs. We exploit this by grouping all tiles (ray tracers, routers, and generator) into a single concurrent compute step that execute parallel, then alternating compute and exchange programs. The resulting IPU program consists of the following subprograms or kernels:
@@ 5.1 Reconstruction quality
Fig. 8 compares our IPU implementation to the Radiant Foam ground truth on the Garden scene. Despite running entirely from on-chip SRAM with simplified per-cell color (constant RGB instead of spherical harmonics), the IPU renderer preserves fine detail with only minor view-dependent effects differences.
@@
Fig. 9 visualizes the same comparison: full-, mixed-, and half-precision payloads produce outputs that are visually indistinguishable.
@@
[…]
@@ 5.2 Scene partitioning and on-chip memory footprint
Figure 11: K-d tree partitioning of Bicycle scene into 1024 clusters (scene partitions). Each cluster corresponds to one leaf of the k-d tree and becomes a scene partition on a ray-tracer tile. For each cluster, the four subplots report the number of local cells ($N_{\text{local}}$), neighbor cells ($N_{\text{neighbor}}$), adjacency entries ($N_{\text{adjacency}}$), and neighboring clusters, respectively.
@@ 5.3 Throughput and batching
Fusing all subprograms( Sec. 4.8) into a single device-side loop shifts the majority of runtime from host engine launches and global barriers into tile-local compute-set execution (Detailed comparison can be found inAppendix C). We further accelerate execution by tuning the repeat count $RC$, the number of compute–exchange iterations executed on device before reading back the output image and updating the camera.
@@
Table 3: Row scanning at $640{\times}480$ with a 200-iteration drain after injection, rendering at approximately 1 frame per second.
@@ 5.4 Routing path length and ray lifetime
(b) Histogram of router and tracer hops per pixel until termination for the Bicycle scene.
Figure 14: Bicycle scene path statistics. The hop map (a) reveals where rays repeatedly cross partition boundaries, while the histogram (b) summarizes the distribution of router and tracer hops over all pixels, with average values of 75.6 router hops and 16.7 tracer hops per ray.
@@ 5.5 Router capacity and ray payload precision
After program memory allocation, each router tile is left with roughly $\approx 576$ kB of SRAM to ten buffers (five inputs, five outputs), so each buffer holds about 57.6 kB of rays. Router capacity therefore depends directly on the ray payload size from  Sec. 4.4. Tab. 4 reports the implied per-buffer limits for different layouts.
@@ 5.6 Limitations and artifacts
Figure 16: Ray-origin mismatch artifact in the Playroom scene when the camera pose changes mid-batch. Rays launched under the old pose later reconstruct their origins using the updated pose, causing misaligned directions and resulting in black or white speckle.
@@ Appendix A Radiant Foam scene representation
Intersecting these half-spaces over all neighbors reconstructs the full Voronoi cell of $p_{c}$ (Figure LABEL:fig:voronoi_reconstruction3).
Appendix B Scene partitioning: K-d tree vs. Octree
@@ B.1 Octree with per-partition point cap $\leq$ 5000 points
Limiting points per partition greatly increases the number of partitions (>>3500), exceeding the available tiles and making a one-tile-per-partition mapping infeasible as shown in Fig. 18.
@@
B.2 Octree with Fixed Partition Count (1024 Partitions)
Forcing 1024 partitions yields highly unbalanced partitions: some partitions exceed $2\times 10^{4}$ local cells and $3\times 10^{5}$ adjacency entries, surpassing 624 kB SRAM per tile even before queues and workspaces as shown in Fig. 19.
@@ D.1 Bicycle scene
(a) Rendered RGB view of the Garden scene.
@@ Appendix F Additional qualitative results
Figures 25 and 26 show paired RGB renderings and depth maps for six scenes. Depth is computed at the 50% transmittance quantile, producing smooth, geometry-aligned results.
(a) Bicycle (RGB)
(b) Bicycle (Depth)
(c) Garden (RGB)
(d) Garden (Depth)
(e) Stump (RGB)
(f) Stump (Depth)
[…]
(a) Bonsai (RGB)
(b) Bonsai (Depth)
(c) Kitchen (RGB)
(d) Kitchen (Depth)
(e) Counter (RGB)
(f) Counter (Depth)
[…]